# Adaptive Joint Estimation of Temporal Vertex and Edge Signals

### v2

@@ Abstract
The adaptive estimation of coexisting temporal vertex (node) and edge signals on graphs is a critical task when a change in edge signals influences the temporal dynamics of the vertex signals. However, the current Graph Signal Processing algorithms mostly consider only the signals existing on the graph vertices and have neglected the fact that signals can reside on the edges. We propose an Adaptive Joint Vertex-Edge Estimation (AJVEE) algorithm for jointly estimating time-varying vertex and edge signals through a time-varying regression, incorporating both vertex signal filtering and edge signal filtering. Accompanying AJVEE is a newly proposed Adaptive Least Mean Square procedure based on the Hodge Laplacian (ALMS-Hodge), which is inspired by classical adaptive filters combining simplicial filtering and simplicial regression. AJVEE is able to operate jointly on the vertices and edges by merging two ALMS-Hodge algorithms specified on the vertices and edges into a unified formulation. A more generalized case extending AJVEE beyond the vertices and edges is being discussed. Experimenting on real-world traffic networks and population mobility networks, we have confirmed that our proposed AJVEE algorithm could accurately and jointly track time-varying vertex and edge signals on graphs.
@@ I Introduction
Graphs have recently become a popular and impactful research topic, owing to the effective representation ability at representing interacting multivariate signals. Four major challenges impede the effectiveness of classical signal processing algorithms on multivariate time-varying signals that reside on graph-like topological structures: the representation of topological irregularity, the estimation of unknown data based on known data, the removal of noise present in the data, and the extraction of the cross-space-time time-variation. In graph signal processing (GSP), graphs have demonstrated their capacity to represent information in various real-world scenarios and disciplines of studies. Graphs have been studied extensively and attracted considerable attention due to their irregular structure, as well as their effectiveness in representing interactions among data [1, 2, 3]. Measurements taken by multiple sensors could be used to construct sensor graphs based on the locations of the sensors, with the data features embedded on the vertices. Examples of sensor graphs include recordings of 5G signal reception strength graph [4], air-quality graph [5], and nationwide temperature graph [6]. In bioinformatics, biological networks could model the time-varying gene interactions [7], functional brain imaging [8], and biomedicine [9]. Graphical models are proposed in finance to reflect the interplay of the stock market from the interaction among stocks [10].
Signals in reality often contain time dimension information; when processing the data on the graph vertices that evolve with time, one could rely on the graph adaptive filters, which is a combination of classical adaptive filters with the graph shift operations. The first GSP adaptive algorithm proposed was the graph adaptive least mean squares (GLMS) algorithm and graph signals are estimated in an online fashion using $l_{2}$-norm optimization with the assumption of Gaussianity on the noise of the signal [4]. Extensions of the GLMS have enhanced the performances of adaptive GSP filters at improved convergence speed [6, 11] and increased robustness under impulsive noise [11, 12, 13, 14]. A graph diffusion based on the adaptive GLMS was proposed in [15] and another graph adaptive filter based on GLMS that uses a sign update strategy was recently proposed in [16]. Kernel-based time-varying graph signal reconstruction method was also proposed using space-time models [17]. Another class of online algorithms that processes time-varying nodal signals is to combine time-series analysis techniques with GSP, for example, the graph vector autoregressive model [18], the graph vector autoregressive moving average model [19, 20], and the graph GARCH model [21]. Time-varying vertex signal can also be processed using offline algorithms based on time-vertex algorithms [22, 23].
@@
Despite the success of GSP, it solely considers data on the graph vertices, meaning that higher-order signals, such as signals on graph edges, are not being represented. To give a few examples, traffic data on road networks recorded on the edges [24, 25], gene interaction over time forms a set of time-varying graph edges [7], multi-agent systems use time-varying edge data to model the agent interactions [26]. A citation complex can be viewed as a higher-ordered social network and formed from constructing simplices among co-author collaborations: the number of citations of a 2-author is a signal on the graph edges and the number of citations of a 3-author paper is a signal on a triangle [27]. The spread of COVID-19 among multiple regions can be represented using a series of dynamic graph that records the confirmed cases in each region on the vertex and the population mobility between two regions on the edges. In existing GSP algorithms that process time-varying vertex signals, the interaction among vertices is modeled using fixed parameters at all time instances [21]. Failing to capture the time-varying interaction would lead to an inaccurate estimation. One could attempt to apply GSP to signals on the edges by projecting the edge signals onto the vertices of its vertex-to-edge dual graph (line graph) as a workaround, but as pointed out in [28] and [29], the projected signals do not share the same GSP properties and assumptions, which hinders the performance of the line graph workaround.
@@
The recently emerged Topological Signal Processing (TSP) has been established as an alternative point of view from GSP for heterogeneous and higher-ordered types of signals, overcoming the limitation of GSP by generalizing graphs to simplicial complexes. The signals on the vertices of a graph are considered as signals of the 00 order simplex, while higher order simplices could be used to represent signals on the edges and beyond [30, 31]. In [24], semi-supervised learning for data on the graph edges is done using the Hodge Laplacians and flow conservation. An analogy of spectral wavelets on simplicial complexes called Hodgelets is proposed in [32]. The work of [33] further proposed topological Slepians, which are the signals concentrated on the topological domain and localized on the frequency domain. In [34, 35], finite impulse response filters on simplicial complexes were discussed, and the work of [36] further defined the convolution operation for signals on a simplicial complex. Chebyshev polynomial approximations of spectral filters on simplicial complexes are provided in [36] and [37]. Recently the discussions on online water flow imputation on graph edges were attempted strictly using flow conservation in [38]; the simplicial vector autoregressive (SVAR) was proposed in [39] on the same water flow imputation task. The sampling of graph edges using the line graph transformation is discussed in [40]. Deep learning algorithms making use of simplicial complexes to process data on the edges have been proposed by introducing nonlinearity to TSP. Graph neural network for edge data based on the Hodge Laplacian can be found in [28]. The work of [27] defined a convolutional neural network based on simplicial complexes. The attention mechanism is merged with the Hodge Laplacian in [41] and [42].
@@
The above Hodge Laplacian-based algorithms demonstrated good performance at processing signals on graph edges, but we may encounter additional challenges in more complicated situations. First and foremost, prior TSP and GSP literature primarily focused on signals in a single simplicial order. The consideration of processing signals that reside in multiple orders of simplices urges increased attention to consider the signals that may appear and interact in multiple orders of simplices. In the previously mentioned biological networks, transportation networks, and population mobility networks examples, the graph vertices and edges both contain signals, with the edge signals influencing the behavior of the vertex signals [7, 26, 43]. However, a deeper discussion should be conducted on the processing of time-varying signals beyond the graph vertices. Numerous real situations require online time-varying data models; static algorithms may not perform optimally to capture the dynamic change in the time-varying data. Understanding the dynamic interactions within these networks is a crucial task for facilitating real-time monitoring and predictive decision-making to enhance efficiency and reliability in real-world applications. On a side note, data gathered from the real world is often noisy, meaning that the denoising of simplicial data should be considered when formulating simplicial representations. These challenges comprise addressing topological irregularities, revealing data interactions across multiple dimensions, eliminating data noise, and capturing time-varying patterns. Therefore, it is a pressing necessity for an online algorithm that focuses on time-varying multi-order and multi-variate data on simplicial complexes to bridge the cross-space-time signal gaps.
@@
In this paper, we propose the Adaptive Joint Vertex-Edge Estimation (AJVEE) algorithm as a strategy aimed at addressing the previously mentioned focuses for time-varying signals existing on both the vertices and the edges by considering the dynamic interplay between the signals on the vertices and edges. To the best of our knowledge, this is the first paper to consider the joint online estimation of time-varying signals on both the vertices and the edges. The contributions of this paper are listed as follows:
@@
  • We propose the AJVEE algorithm on graphs to conduct an adaptive estimation of time-varying vertex and edge signals. AJVEE uses a time-varying regression model to capture the interaction of a time-varying vertex signal, with the intensity of the interactions represented by a time-varying signal on the graph edges.
@@
  • An adaptive Least Mean Square procedure based on the Hodge Laplacian (ALMS-Hodge) is derived to facilitate AJVEE. The resulting ALMS-Hodge is suitable for processing time-varying signals on simplicial complexes using TSP. Different from previous adaptive GSP algorithms that operate only on the graph vertices estimation errors, the ALMS-Hodge can be deployed on higher-ordered structures such as the edges, while considering both the aggregated signal and the estimation error.
  • By specifying one ALMS-Hodge on the vertices and another on the edges, then merging them through a time-varying regression incorporating both the node signal shifts and edge signal shifts, AJVEE is able to operate jointly on the vertices and edges. Additionally, by generalizing the ALMS-Hodge to simplicial complexes with higher orders than the vertices and edges, we can formulate and extend AJVEE beyond the vertices and edges.
@@
In Section II, we will provide preliminary knowledge in representing graphs using simplicial complexes, as well as basic signal processing techniques on simplicial complexes. Derivations and analysis of the AJVEE are included in Section III. Section IV gives a more generalized formulation of AJVEE applied to time-varying signals on simplicial complexes that are not restricted simply to the vertices and edges. Experiment results on both synthetic data and real data are discussed in Section V. Section VI concludes the paper with a brief discussion of potential future work.
@@ II-A Graph representation
To represent the relationship between two simplices of order $k-1$ and $k$ in a simplicial complex, the incidence matrix $\mathbf{B}_{k}\in\mathbb{R}^{N_{k-1}\times N_{k}}$ is used where the rows of $\mathbf{B}_{k}$ correspond to the $N_{k}$ elements in $\mathcal{X}_{k-1}$ and the columns of $\mathbf{B}_{k}$ correspond to the $N_{k}$ elements in $\mathcal{X}_{k}$. The values assignment of $\mathbf{B}_{x}$ are as follows: if simplex $x_{k-1}(i)$ in $\mathcal{X}_{k-1}$ is adjacent to a simplex $x_{k}(j)$ in $\mathcal{X}_{k}$, then the $ij^{th}$ element of $\mathbf{B}_{k}$ will have magnitude $1$, with the sign determined by the predefined orientation in the $k$-simplex. We should emphasize that the orientation, defined by the ordering of the simplicial set, is not the same as the direction, especially when $k=1$ for the edges [44]. Notice that the incidence matrix $\mathbf{B}_{k}$ is a boundary operator because $\mathbf{B}_{k}$ represents how $k-1$-simplices bounds their upper adjacent $k$-simplices. Similarly, we can define the dual of the boundary operator by the transpose of the incidence matrix $\mathbf{B}_{k}^{T}$, which is a coboundary operator.
@@ II-B Signal Processing on graphs and simplicial complexes
With the SFT defined, we can apply simplicial filters $\mathbf{\Sigma}$ to each frequency component in $\mathbf{L}_{k}$ to conduct spectral domain operations on simplicial complexes. We could define several spectral TSP filters and linearly combine them to obtain the desired frequency response. A simple bandlimited filter in $\mathcal{X}_{k}$ based on a frequency set $\mathcal{F}_{k}$ is constructed by a diagonal matrix $\mathbf{\Sigma}_{\mathcal{F},k}$, where a $1$ on the $i^{th}$ diagonal element indicates that $\lambda_{i}\in\mathcal{F}_{k}$ and 00 otherwise [30]. The following procedure is a simple yet complete bandlimited spectral filtering of $\boldsymbol{x}_{k}$:
@@
If the signal $\boldsymbol{x}_{k}$ is bandlimited with frequencies $\mathcal{F}_{k}$, then $\boldsymbol{x}_{k}=\mathbf{U}_{k}\mathbf{\Sigma}_{\mathcal{F},k}\mathbf{U}_{k}^{T}\boldsymbol{x}_{k}$. To simplify notation and make use of the sparsity gained from eliminating some frequency components, we will set $\mathbf{U}_{k,\mathcal{F}}=$ support$(\mathbf{U}_{k}\mathbf{\Sigma}_{\mathcal{F},k})$, where support() is the operation of dropping the column with all zero elements. Notice that the filter $\mathbf{\Sigma}_{\mathcal{F},k}$ is idempotent and self-adjoint, so with the simplified notation we have $\mathbf{U}_{k,\mathcal{F}}\mathbf{U}_{k,\mathcal{F}}^{T}=\mathbf{U}_{k}\mathbf{\Sigma}_{\mathcal{F},k}\mathbf{U}_{k}^{T}$, and $\mathbf{U}_{k,\mathcal{F}}^{T}\mathbf{U}_{k,\mathcal{F}}=\mathbf{I}$. According to the TSP sampling theory for signals defined on $\mathcal{K}$, partial observation is modeled using a sampling operation on $\mathbf{D}_{k,\mathcal{S}}$ [30]. The sampling set $\mathcal{S}_{k}$ defines the diagonal matrix masking operation, where the sampling matrix $\mathbf{D}_{k,\mathcal{S}}=\text{diag}(d_{k,1}...d_{k,N_{k}})$, where $d_{k,i}=1$ if $x_{k,i}\in\mathcal{S}_{k}$ and 00 otherwise [30]. The sampling matrix $\mathbf{D}_{k,\mathcal{S}}$ is idempotent and self-adjoint as well.

### v3

@@ Abstract
The adaptive estimation of coexisting temporal vertex (node) and edge signals on graphs is a critical task when a change in edge signals influences the temporal dynamics of the vertex signals. However, the current Graph Signal Processing algorithms mostly consider only the signals existing on the graph vertices and have neglected the fact that signals can reside on the edges. We propose an Adaptive Joint Vertex-Edge Estimation (AJVEE) algorithm for jointly estimating time-varying vertex and edge signals through a time-varying regression, incorporating both vertex signal filtering and edge signal filtering. Accompanying AJVEE is a newly proposed Adaptive Least Mean Square procedure based on the Hodge Laplacian (ALMS-Hodge), which is inspired by classical adaptive filters combining simplicial filtering and simplicial regression. AJVEE is able to operate jointly on the vertices and edges by merging two ALMS-Hodge specified on the vertices and edges into a unified formulation. A more generalized case extending AJVEE beyond the vertices and edges is being discussed. Experimenting on real-world traffic networks and population mobility networks, we have confirmed that our proposed AJVEE algorithm could accurately and jointly track time-varying vertex and edge signals on graphs.
@@ I Introduction
Graphs have recently become a popular and impactful research topic due to their effective ability to represent interacting multivariate signals. However, several challenges must be addressed to enhance the effectiveness of signal processing algorithms on multivariate time-varying signals that reside on graph-like topological structures. These challenges include representing topological irregularity, estimating unknown data from known data, removing noise present in the data, and extracting cross-space-time variations. In graph signal processing (GSP), graphs have demonstrated their capacity to represent information in various real-world scenarios and disciplines of studies. Graphs have been studied extensively and attracted considerable attention due to their irregular structure, as well as their effectiveness in representing interactions among data [1, 2, 3]. Measurements taken by multiple sensors could be used to construct sensor graphs based on the locations of the sensors, with the data features embedded on the vertices. Examples of sensor graphs include recordings of 5G signal reception strength graph [4], air-quality graph [5], and nationwide temperature graph [6]. In bioinformatics, biological networks can model time-varying gene interactions. [7], functional brain imaging [8], and biomedicine [9]. Graphical models are proposed in finance to reflect the interplay of the stock market from the interaction among stocks [10].
@@
Despite the success of GSP, GSP solely considers data on the graph vertices, meaning that higher-order signals, such as signals on graph edges, are not being represented. To provide a few examples, traffic data on road networks recorded on the edges [11, 12], gene interaction over time forms a set of time-varying graph edges [7], multi-agent systems use time-varying edge data to model the agent interactions [13]. A citation complex can be viewed as a higher-ordered social network and formed from constructing simplices among co-author collaborations: the number of citations of a 2-author is a signal on the graph edges and the number of citations of a 3-author paper is a signal on a triangle [14]. The spread of COVID-19 across multiple regions can be represented using a series of dynamic graphs, where the vertices record the confirmed cases in each region, and the edges represent population mobility between regions. In existing GSP algorithms that process time-varying vertex signals, the interaction among vertices is modeled using fixed parameters at all time instances [15]. Failing to capture the time-varying interaction would lead to an inaccurate estimation.
@@
The recently emerged Topological Signal Processing (TSP) has been established as an alternative point of view from GSP for higher-ordered signals, overcoming the limitation of GSP by generalizing graphs as simplicial complexes. The signals on the vertices of a graph are considered as signals of the 00 order simplex, while higher order simplices could be used to represent signals on the edges and beyond [16, 17]. In [11], semi-supervised learning for data on the graph edges is done using the Hodge Laplacians and flow conservation. An analogy of spectral wavelets on simplicial complexes called Hodgelets is proposed in [18]. The work of [19] further proposed topological Slepians, which are the signals concentrated on the topological domain and localized on the frequency domain. In [20, 21], finite impulse response filters on simplicial complexes were discussed, and the work of [22] further defined the convolution operation for signals on a simplicial complex. Recently, the proposal of using a Dirac operator allows the merging of fixed time-invariant signals from different orders of simplex [23]. Chebyshev polynomial approximations of spectral filters on simplicial complexes are provided in [22] and [24]. The sampling of graph edges using the line graph transformation is discussed in [25].
The above TSP methods demonstrated good performance at processing signals on graph edges, but we may encounter additional challenges in realistic situations. Signals, in reality, often contain information about time dimension; when processing data on graph vertices that evolve with time, one could rely on graph adaptive filters, a combination of classical adaptive filters with graph shift operations. In GSP, the graph adaptive least mean squares (GLMS) algorithm was proposed for online estimation of vertex signals through $l_{2}$-norm optimization, assuming Gaussian noise and bandlimitedness of signal [4]. Extensions of GLMS have improved its performance in various aspects. For instance, the normalized GLMS (GNLMS) incorporates spectral domain normalization in the bandlimited filter, enhancing convergence speed compared to GLMS [6]. Other algorithms such as the Graph Least Mean $p^{th}$ (GLMP) [26], normalized GLMP [27], and Graph Alternating Minimization [28] algorithms have increased robustness against impulsive noise. The adaptive Graph-Sign (G-Sign) reduces computational complexity and increases robustness by employing an update function based on $l_{1}$ norm optimization [29, 30]. Another class of algorithms that processes time-varying vertex signals combines time-series analysis techniques such as Vector Autoregressive (VAR), Vector Autoregressive Moving Average (VARMA), and generalized autoregressive conditionally heteroscedastic (GARCH) models GSP, resulting in the graph VAR [31], graph VARMA [32, 33], and graph GARCH [15]. Recently, the discussions of time-varying signals beyond the graph vertices have emerged: online water flow imputation on graph edges was attempted strictly using flow conservation in [34]; the simplicial VAR was proposed in [35] on the same water flow imputation task. The major drawback of implementing time-series analysis techniques on topological structures is that the multivariate nature of topological signals increases model complexity and poses challenges in acquiring accurate models. Research on the online estimation of signals on simplicial complexes is limited, particularly for the joint task of spatial signal imputation and temporal adaptive predictions.
@@
Prior TSP and GSP methods have another limitation of primarily focusing on signals in a specific simplicial order. In other words, the task to be solved is only on the vertices or the edges, but rarely both, not to mention that the majority of previous GSP and TSP methods are targeted for time-invariant scenarios. The need to process signals residing in multiple orders of simplices has drawn increased attention to the interactions and appearances of signals across these different simplicial orders. In biological networks, transportation networks, and population mobility networks examples, the graph vertices and edges both contain signals, with the edge signals influencing the behavior of the vertex signals [7, 13, 36]. Understanding the dynamic interactions within these networks and signals from different simplicial orders is a crucial task for facilitating online tracking and predictive decision-making algorithms to enhance efficiency and reliability in real-world applications. On a side note, data gathered from the real world is often noisy, meaning that the denoising of simplicial data should be considered when formulating simplicial representations. These challenges comprise addressing topological irregularities, revealing data interactions across multiple dimensions, eliminating data noise, and capturing time-varying patterns. Therefore, it is a pressing necessity for an online algorithm that focuses on time-varying multi-order and multi-variate data on simplicial complexes to bridge the cross-space-time signal gaps.
@@
In this paper, we propose the Adaptive Joint Vertex-Edge Estimation (AJVEE) algorithm as a strategy aimed at addressing the previously mentioned focuses for time-varying signals existing on both the vertices and the edges by considering the dynamic interplay between the signals on the vertices and edges. To the best of our knowledge, this is the first paper to consider the joint online estimation of time-varying signals on both the vertices and the edges. The major contributions of this paper are listed as follows:
@@
  • The AJVEE framework is proposed for adaptive estimation of time-varying vertex and edge signals on graphs by integrating two separate online estimations for vertices and edges and then into a single time-varying regression model. AJVEE utilizes a time-varying regression model to capture the interaction between vertex signals, with the intensity of these interactions represented by edge signals. AJVEE can be further generalized to operate beyond just vertices and edges, allowing for more complex signal interactions.
@@
  • An adaptive Least Mean Square procedure based on the Hodge Laplacian (ALMS-Hodge) is derived to support AJVEE. Unlike previous adaptive GSP algorithms that only address vertex estimation errors, ALMS-Hodge is generalized to higher-order structures such as edges and triangles. ALMS-Hodge relies upon both aggregated signals and estimation errors to conduct missing signal imputation in the spatial domain and adaptive predictions in the temporal domain.
@@
In Section II, we will provide preliminary knowledge in representing graphs using simplicial complexes, as well as basic signal processing techniques on simplicial complexes. The ALMS-Hodge is introduced in Section III. Derivations and analysis of the AJVEE are included in Section IV; we also provide a more generalized formulation of AJVEE applied to time-varying signals on simplicial complexes that are not restricted simply to the vertices and edges. Experiment results on both synthetic data and real data are discussed in Section V. Section VI concludes the paper with a brief discussion of potential future work. Table I summarizes some frequent symbols seen in this paper.
TABLE I: Symbols and their meanings
Symbol Meaning
$k$ Order of a simplex or simplicial complex
$\mathcal{X}_{k}$ $k$-simplex (vertex: $k$ = 0, edge: $k$ = 1)
$\mathcal{K}$ simplicial complex
$\mathcal{G}$ graph
$t$ time index
$\boldsymbol{x}_{k}[t]\in\mathbb{R}^{N_{k}}$ $k$-simplicial signal of time $t$
$N_{k}$ Cardinality of $\boldsymbol{x}_{k}[t]$
$\mathbf{L}_{k}\in\mathbb{R}^{N_{k}\times N_{k}}$ Hodge Laplacian matrix
$\mathbf{D}_{k,\mathcal{S}}\in\mathbb{R}^{N_{k}\times N_{k}}$ Sampling (Masking) Matrix
$\mathbf{U}_{k}\in\mathbb{R}^{N_{k}\times N_{k}}$ Eigenvector matrix of $\mathbf{L}_{k}$
$\mathcal{F}_{k}$ The frequency set of the bandlimited filter
$\mathbf{U}_{k,\mathcal{F}}\in\mathbb{R}^{N_{k}\times|\mathcal{F}_{k}|}$ Bandlimited eigenvector matrix of $\mathbf{L}_{k}$
$\mathbf{B}_{k}\in\mathbb{R}^{N_{k-1}\times N_{k}}$ Incidence matrix
$\mathbf{I}$ Identity matrix
$\mathbf{\Sigma}_{k}\in\mathbb{R}^{N_{k}\times N_{k}}$ Simplicial spectral filter
$\hat{\boldsymbol{x}}_{k}[t]\in\mathbb{R}^{N_{k}}$ Estimated $k$-simplicial signal of time $t$
$\hat{\boldsymbol{s}}_{k}[t]\in\mathbb{R}^{N}$ SFT of $\hat{\boldsymbol{x}}_{k}[t]$
$\boldsymbol{\eta}_{k}[t]\in\mathbb{R}^{N_{k}}$ Noise of the $k$-simplicial signal
${\boldsymbol{y}}_{k}[t]\in\mathbb{R}^{N_{k}}$ Noisy observation of ${\boldsymbol{x}}_{k}[t]$
$\mu_{k}$ Update step size
$r_{k}$ Aggregation weight
@@ II-A Graph representation
To represent the relationship between two simplices of order $k-1$ and $k$ in a simplicial complex, the incidence matrix $\mathbf{B}_{k}\in\mathbb{R}^{N_{k-1}\times N_{k}}$ is used where the rows of $\mathbf{B}_{k}$ correspond to the $N_{k}$ elements in $\mathcal{X}_{k-1}$ and the columns of $\mathbf{B}_{k}$ correspond to the $N_{k}$ elements in $\mathcal{X}_{k}$. The values assignment of $\mathbf{B}_{k}$ are as follows: if simplex $x_{k-1}(i)$ in $\mathcal{X}_{k-1}$ is adjacent to a simplex $x_{k}(j)$ in $\mathcal{X}_{k}$, then the $ij^{th}$ element of $\mathbf{B}_{k}$ will have magnitude $1$, with the sign determined by the predefined orientation in the $k$-simplex. We should emphasize that the orientation, defined by the ordering of the simplicial set, is not the same as the direction, especially when $k=1$ for the edges [37]. The incidence matrix $\mathbf{B}_{k}$ is a boundary operator because $\mathbf{B}_{k}$ represents how $k-1$-simplices bounds their upper adjacent $k$-simplices. Similarly, we can define the dual of the boundary operator by the transpose of the incidence matrix $\mathbf{B}_{k}^{T}$, which is a coboundary operator [16].
@@ II-B Signal Processing on graphs and simplicial complexes
With the SFT defined, we can apply predefined simplicial filters $\mathbf{\Sigma}$ to each frequency component in $\mathbf{L}_{k}$ to conduct spectral domain operations on simplicial complexes. A simple bandlimited filter in $\mathcal{X}_{k}$ based on a frequency set $\mathcal{F}_{k}$ is constructed by a diagonal matrix $\mathbf{\Sigma}_{\mathcal{F},k}$, where a $1$ on the $i^{th}$ diagonal element indicates that $\lambda_{i}\in\mathcal{F}_{k}$ and 00 otherwise [16]. We can define several spectral TSP filters and linearly combine them to obtain the desired frequency response. The following procedure is a simple yet complete bandlimited spectral filtering of $\boldsymbol{x}_{k}$:
@@
If the signal $\boldsymbol{x}_{k}$ is bandlimited with frequencies $\mathcal{F}_{k}$, then $\boldsymbol{x}_{k}=\mathbf{U}_{k}\mathbf{\Sigma}_{\mathcal{F},k}\mathbf{U}_{k}^{T}\boldsymbol{x}_{k}$. To simplify notation and make use of the sparsity gained from eliminating some frequency components, we will set $\mathbf{U}_{k,\mathcal{F}}=$ support$(\mathbf{U}_{k}\mathbf{\Sigma}_{\mathcal{F},k})$, where support() is the operation of dropping the column with all zero elements. Notice that the filter $\mathbf{\Sigma}_{\mathcal{F},k}$ is idempotent and self-adjoint, so with the simplified notation we have
$\mathbf{U}_{k,\mathcal{F}}\mathbf{U}_{k,\mathcal{F}}^{T}=\mathbf{U}_{k}\mathbf{\Sigma}_{\mathcal{F},k}\mathbf{U}_{k}^{T}$ (4)
and
$\mathbf{U}_{k,\mathcal{F}}^{T}\mathbf{U}_{k,\mathcal{F}}=\mathbf{I}.$ (5)
The filtering operation in (3) can be done using a spectral approach or a spatial approach [24]. To achieve similar effects as (3) in the spatial domain, a predefined bandlimited filter is approximated using a series of shifted Chebyshev polynomials, and the spatial relationship is represented by the Hodge Laplacian. Assuming that spatial operations are conducted in the $k$-simplicies of simplicial complex $\mathcal{K}$, the shifted Chebyshev polynomial $T_{p}(\mathbf{L}_{k})$ with order $P$ is defined as
$T_{p}(\mathbf{L}_{k})=\begin{cases}1,&\text{if }p=0,\\ \frac{2\mathbf{L}_{k}-\lambda_{\text{max}}}{\lambda_{\text{max}}},&\text{if }p=1,\\ \frac{4\mathbf{L}_{k}-2\lambda_{\text{max}}}{\lambda_{\text{max}}}T_{p-1}(\mathbf{L}_{k})-T_{p-2}(\mathbf{L}_{k}),&\text{if }p\geq 2.\end{cases}$ (6)
Using (6), the approximation of $\mathbf{U}_{k}\mathbf{\Sigma}_{\mathcal{F},k}\mathbf{U}_{k}^{T}$ in (8) is expressed as
$\mathbf{U}_{k}\mathbf{\Sigma}_{\mathcal{F},k}\mathbf{U}_{k}^{T}\approx{\theta_{0}}+\sum_{p=1}^{P}\theta_{p}T_{p}(\mathbf{L}_{k}),$ (7)
where $\theta_{p}$ is the coefficient of the Chebyshev polynomial.
A special case of the bandlimited filter is the ideal low-pass filter, which is based on the (spatial) smoothness assumption of the signal [25]. The ideal low-pass filter $\mathbf{\Sigma}_{k,\mathcal{F}_{lp}}$ with cutoff frequency $\lambda_{k,c}$ can be defined using the frequency set $\mathcal{F}_{lp}$ and can be expressed as
$\mathbf{\Sigma}_{k,\mathcal{F}_{lp}}=\text{diag}(\boldsymbol{\sigma}_{k,\mathcal{F}_{lp}}),$ (8)
where $\lambda_{k,i}\in\mathcal{F}_{lp}\;|\;i\leq c,\quad\lambda_{k,i}\notin\mathcal{F}_{lp}\;|\;i>c$, and $1<c<N_{k}$ is the index of the cutoff frequency $\lambda_{k,c}$. This means that the value assignment of the $i^{th}$ element of $\boldsymbol{\sigma}_{k,\mathcal{F}_{lp}}$ follows ${\sigma}_{k,\mathcal{F}_{lp},i}=\mathbb{I}(i\leq c)$.
According to the TSP sampling theory for signals defined on $\mathcal{K}$, partial observation is modeled using a sampling operation on $\mathbf{D}_{k,\mathcal{S}}$ [16]. The sampling set $\mathcal{S}_{k}$ defines the diagonal matrix masking operation, where the sampling matrix $\mathbf{D}_{k,\mathcal{S}}=\text{diag}(d_{k,1}...d_{k,N_{k}})$, where $d_{k,i}=1$ if $x_{k,i}\in\mathcal{S}_{k}$ and 00 otherwise [16]. The sampling matrix $\mathbf{D}_{k,\mathcal{S}}$ is a diagonal matrix that is both idempotent and self-adjoint. When signal observations are missing due to various conditions or criteria other than by sampling, the matrix $\mathbf{D}_{k,\mathcal{S}}$ still denotes the observation mask.